# Equivalence of Privacy and Stability with Generalization Guarantees in Quantum Learning

### v1

@@ Abstract
We present a unified information-theoretic framework to analyze the generalization performance of differentially private (DP) quantum learning algorithms. By leveraging the connection between privacy and algorithmic stability, we establish that $(\varepsilon,\delta)$-Quantum Differential Privacy (QDP) imposes a strong constraint on the mutual information between the training data and the algorithm’s output. We derive a rigorous, mechanism-agnostic upper bound on this mutual information for learning algorithms satisfying a 1-neighbor privacy constraint. Furthermore, we connect this stability guarantee to generalization, proving that the expected generalization error of any $(\varepsilon,\delta)$-QDP learning algorithm is bounded by the square root of the privacy-induced stability term. Finally, we extend our framework to the setting of an untrusted Data Processor, introducing the concept of Information-Theoretic Admissibility (ITA) to characterize the fundamental limits of privacy in scenarios where the learning map itself must remain oblivious to the specific dataset instance.
@@ I-B Our Framework and Main Contributions
First step: Stability implies Generalizability. We begin by addressing the Investigator’s primary goal: to extract useful information that generalizes to unseen data. In the classical setting, the relationship between algorithmic stability and generalization has been extensively studied. Bousquet and Elisseeff [3] established a foundational connection, showing that if a learning algorithm is uniformly stable—meaning its output is insensitive to small perturbations in the training set—it guarantees tight bounds on the generalization error. This ensures that the empirical error observed on the training data is a reliable proxy for the true error on future data. Extending this classical intuition to the quantum domain, we prove that stable quantum algorithms possess similar strong generalization capabilities. By assuming a classical-quantum sub-Gaussian property for the loss function, we derive a bound on the expected generalization error in terms of the mutual information between the training data and the output. This result confirms that if an algorithm does not depend too heavily on any single quantum data point, it prevents overfitting.
@@
Second step: Privacy implies Generalizability via Stability. Having established that stability is the key to generalization, we turn to the Respondent’s perspective: how do we ensure this stability in a rigorous manner? The answer lies in Differential Privacy. From the Respondent’s perspective, privacy necessitates that the Investigator cannot discern the inclusion or exclusion of any individual’s data. Differential privacy formalizes this requirement as a strong, mathematically precise form of stability. In the classical case, Dwork et al. [10] explicitly linked this to learning, demonstrating that differential privacy automatically ensures stability, and consequently, provides generalization guarantees. The main contribution of this manuscript is to generalize these results to the quantum learning scenario for the more general $(\varepsilon,\delta)$-DP setting. We propose a framework for a quantum learning algorithm which is "1-neighbor" $(\varepsilon,\delta)$-DP and study its stability by obtaining a rigorous upper-bound on the mutual information between the training data and the quantum output. By combining this with the result of the first step, we formally demonstrate that privacy is not merely a constraint, but a sufficient condition for generalization in quantum learning.
Third step: Dishonest Third Party and Information-Theoretic Admissibility. While the previous steps assume a trusted Data Processor who performs the learning algorithm, we also address the scenario where the third party, the Data Processor, cannot be trusted. In this untrusted Data Processor scenario, the learning algorithm must be independent of the specific training data instance to prevent the Data Processor from learning sensitive information. To rigorously characterize this setting, we introduce the concept of Information-Theoretic Admissibility (ITA). A learning algorithm is considered ITA with respect to a set of states if no other learning algorithm is strictly more informative—meaning no other algorithm allows the Investigator to recover the original information state via a CP-TP map.
If an algorithm is not ITA, the Investigator could potentially construct a "better" algorithm to extract more information, rendering the privacy definitions ineffective. We show that in the quantum setting, unlike the classical case where ITA implies full recoverability of information (meaning privacy is impossible under ITA), there exist non-trivial ITA algorithms where the security condition remains meaningful. Thus, the quantum setting provides a genuine advantage, validating the necessity of our privacy definitions even under the strict ITA condition.
I-C Organization of this Manuscript
The remainder of this manuscript is organized as follows. Section II establishes the requisite mathematical notation. In Section III, we present the general framework for quantum learning and formally define information-theoretic stability. Crucially, we prove that stability implies generalization (Step 2) by deriving an error bound based on a Classical-Quantum Sub-Gaussianity assumption. Section IV applies this framework to Differentially Private (DP) algorithms. We define the $1$-neighbor $(\varepsilon,\delta)$-DP condition and prove that privacy implies stability (Step 1), thereby establishing generalization guarantees for private quantum algorithms. We also compare our results with existing literature. Section V extends the framework to the untrusted Data Processor scenario, introducing Information-Theoretic Admissibility (ITA) to characterize privacy limits (Step 3). Finally, we conclude in Section VI.
@@ II Notations
$\mathcal{D}(\mathcal{H})$ denotes the set of density operators on a finite-dimensional Hilbert space $\mathcal{H}$. The $\log$ function used in the discussions below is to the base $e$.
@@
Let $\mathcal{T}$ be a finite alphabet with $\left|\mathcal{T}\right|=d$. The type of a sequence $t\in\mathcal{T}^{n}$ is the frequency vector ${\bf f}=(f_{1},\ldots,f_{d})\in\mathbb{N}_{0}^{d}$ such that $\sum_{i}f_{i}=n$, where $f_{i}$ counts the occurrences of the $i$-th symbol of $\mathcal{T}$ in $t$. We denote the set of all such types by $T_{d}^{n}$. The type class $T_{\bf f}\subset\mathcal{T}^{n}$ is the set of all sequences sharing the same type $\bf f$.
@@ III-A Learning Setup and Data Encoding
The process begins with the Respondent. The input consists of a classical dataset $s:=(z_{1},\cdots,z_{n})\in\mathcal{S}$ (where $n\in\mathbb{N},\left|\mathcal{Z}\right|>0$) collected from $n$ individual entries of the Respondent. Each Respondent contributes a corresponding quantum state $\rho_{z_{i}}\in\mathcal{D}(\mathcal{H}^{\bm{\mathfrak{Tr}}}\times\mathcal{H}^{\bm{\mathfrak{Te}}})$. Thus, the aggregate input state is $\rho_{s}:=\rho_{z_{1}}\otimes\cdots\otimes\rho_{z_{n}}$. This state consists of two parts: a training part over system $\bm{\mathfrak{Tr}}$ (accessible to the Data Processor) and a testing part over system $\bm{\mathfrak{Te}}$ (used for evaluation), which may be entangled. Each data point $z_{i}:=(x_{i},y_{i})$ maps an input $x_{i}$ to a label $y_{i}$ via an unknown concept function $f$.
The Data Processor receives this input, formalized as a classical-quantum state $\sum_{s}P_{S}(s)|s\rangle\langle s|\otimes\rho_{s}$, and executes a learning algorithm. We model this algorithm as a collection of quantum instruments [8] $\mathcal{N}:=\{\mathcal{N}^{(s)}:\mathcal{H}^{\bm{\mathfrak{Tr}}}\to\mathcal{H}^{B}\}_{s\in\mathcal{S}}$ which acts only on the input training system $\bm{\mathfrak{Tr}}$. The Processor produces an output on system $B$. Thus, the joint state between the input classical data system $S$, the input quantum testing system $\bm{\mathfrak{Te}}$, and the Data Processor’s output system $B$ is given by:
@@ III-C Stability Implies Generalizability For Quantum Learning Algorithms
Definition 2 (Expected Empirical Loss [6, Definition $11$]).
@@
$\hat{L}_{\rho}(\mathcal{N})$ $:=\mathbb{E}_{(S,W)\sim{P}^{\mathcal{N}}_{SW}}[\mathrm{Tr}[L(S,W)(\sigma^{\mathcal{N}}_{S,W})^{\bm{\mathfrak{Te}}B^{\prime}}]]$
$=\mathrm{Tr}[L^{S\bm{\mathfrak{Te}}WB^{\prime}}\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}].$
Definition 3 (Expected True Loss [19, Definition $19$]).
@@
$L_{\rho}(\mathcal{N})$ $:=\mathbb{E}_{(\overline{S},\overline{W})\sim P_{S}\times{P}^{\mathcal{N}}_{{W}}}\left[\mathrm{Tr}\left[L(\overline{S},\overline{W})\left(\rho^{\bm{\mathfrak{Te}}}_{\overline{S}}\otimes(\sigma^{\mathcal{N}}_{\overline{W}})^{B^{\prime}}\right)\right]\right]$
$=\mathrm{Tr}[L^{S\bm{\mathfrak{Te}}WB^{\prime}}(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})].$
@@
We adopt the definition of expected true loss as proposed in [19] and not that of [6]. The authors in [19] give a rigorous justification for Definition 3 and argue that the definition proposed by [6, Definition $12$] is not a correct definition for the expected true loss.
@@
$\overline{\text{{gen}}}_{\rho}(\mathcal{N})$ $:=\left|\hat{L}_{\rho}(\mathcal{N})-L_{\rho}(\mathcal{N}))\right|$
$=\left|\mathrm{Tr}[L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}]-\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\right]\right|.$
@@
Definition 5 (Classical-Quantum Sub-Gaussianity).
For a fixed parameter $\alpha\in(0,\infty),$ the collection $\{L(w,s)\}$ of loss operators is said to be an $\alpha$ sub-Gaussian collection with respect to $\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}:=\sum_{(s,w)\in\mathcal{S}\times\mathcal{W}}P_{S}(s)|s\rangle\langle s|\otimes P^{\mathcal{N}}_{W}(w\mid s)|w\rangle\langle w|\otimes\rho^{\bm{\mathfrak{Te}}}_{s}\otimes(\sigma^{\mathcal{N}}_{w})^{B^{\prime}},$ if for every $\lambda\in\mathbb{R},$ it satisfies,
@@
Definition 5 naturally generalizes classical sub-Gaussianity. In the limit of trivial quantum systems ($\dim(\bm{\mathfrak{Te}})=\dim(B^{\prime})=1$), the operators $\sigma^{\mathcal{N}}_{W}$ and $\mathbb{I}^{B^{\prime}}$ become scalars, reducing $L(S,W)$ to a classical random loss function. Consequently, condition (6) collapses to the standard classical sub-Gaussian inequality
$\mathbb{E}_{(S,W)}[e^{\lambda(L(S,W)-\mathbb{E}[L(S,W)])}]\leq e^{\frac{\lambda^{2}\alpha^{2}}{2}},$
with respect to $P_{S}\times{P}^{\mathcal{N}}_{{W}}$. We now present a theorem bounding the expected generalization error for quantum learning algorithms.
@@
See Appendix D for the proof.
Remark 2.
In [6, Theorem $17$], the authors impose the following two separate point-wise holding sub-Gaussianity requirements,
An important implication which follows from Theorem 1 and Definition 1 is the following corollary.
@@
This corollary establishes that if the Data Processor utilizes a stable algorithm, the generalization error is tightly controlled.
Remark 3.
Unlike the bounds in [6] and [19], our bound on $\overline{\text{{gen}}}_{\rho}(\mathcal{N})$ relies on a single information-theoretic quantity. In scenarios where testing and training data are uncorrelated, the bound in [6] reduces to a purely classical term and therefore it will not account for the quantum system $B^{\prime}$. Therefore, from [6] it is not possible to show that stability implies generalizability. In contrast, Theorem 1 avoids this limitation, validating the definition of expected true loss proposed in [19] as its correct formulation. A justification for the correctness of (3) is also given in [19].
With this correct definition of true loss the bound obtained in [6, Corollary 23] translates to [19, Theorem 1]. In Appendix H, we make a comparison of Theorem 1 with [19, Theorem 1] for the case when $\alpha=\mu=\tau,$ where $\alpha,\mu$ and $\tau$ denote the sub-Gaussianity parameters appearing in (6), (QMGF), and (CMGF), respectively.

### v2

@@ Abstract
We present a unified information-theoretic framework elucidating the interplay between stability, privacy, and the generalization performance of quantum learning algorithms. We establish a bound on the expected generalization error in terms of quantum mutual information and derive a probabilistic upper bound that generalizes the classical result by Esposito et al. (2021). Complementing these findings, we provide a lower bound on the expected true loss relative to the expected empirical loss.
Additionally, we demonstrate that $(\varepsilon,\delta)$-quantum differentially private learning algorithms are stable, thereby ensuring strong generalization guarantees. Finally, we extend our analysis to dishonest learning algorithms, introducing Information-Theoretic Admissibility (ITA) to characterize the fundamental limits of privacy when the learning algorithm is oblivious to specific dataset instances.
@@ I-B Our Framework and Main Contributions
First step: Stability implies Generalizability. We first address the Investigator’s goal of generalization. Building on the classical equivalence between uniform stability and generalization [5], we prove that stable quantum algorithms inherently generalize. Specifically, by assuming a classical-quantum sub-Gaussian property for the loss function, we derive a bound on the expected generalization error in terms of the mutual information between the training data and the output. This confirms that limiting the algorithm’s dependence on individual quantum data points effectively prevents overfitting. Going further, we prove a bound on generalization error in probability using the Sandwiched Rényi divergence [19] to ensure robust performance guarantees. This bound is the quantum version of [13, Corollary $2$]. Finally, we complement these upper bounds with a lower bound on the expected true loss in terms of the empirical loss, establishing a tight bidirectional relationship between observed and actual risk.
@@
Second step: Privacy implies Generalizability via Stability. Next, we address the Respondent’s privacy via Differential Privacy, which acts as a rigorous form of stability. Extending prior results in classical [11] and pure quantum DP [22], our main contribution is to generalize this link to the $(\varepsilon,\delta)$-DP setting. We propose a "1-neighbor" $(\varepsilon,\delta)$-DP framework and derive a rigorous upper bound on the mutual information between the training data and the quantum output. By combining this with our first step, we formally demonstrate that privacy is a sufficient condition for generalization in quantum learning.
Third step: Dishonest Third Party and Information-Theoretic Admissibility. We extend our framework to the untrusted Data Processor scenario, where the learning map must remain independent of specific training instances to prevent leakage to the processor. To rigorously characterize the limits of this setting, we introduce Information-Theoretic Admissibility (ITA). An algorithm is ITA if it is information-theoretically optimal; specifically, no other algorithm exists that is strictly more informative (allowing superior state recovery via a CP-TP map). Non-ITA algorithms render privacy definitions ineffective, as an adversary could employ a superior extraction strategy. Crucially, we demonstrate a fundamental separation: while classical ITA implies full data recoverability (precluding privacy), the quantum setting permits non-trivial ITA algorithms where privacy remains meaningful. This highlights a genuine quantum advantage, validating our security definitions even under the constraint of optimal information extraction.
I-C Our Contributions
Our work comprehensively characterizes the interplay between privacy, stability, and generalization in quantum learning:
  • Expected Generalization: We introduce Classical-Quantum Sub-Gaussianity to bound the expected generalization error by the square root of mutual information (Theorem 1). This unifies classical and quantum fluctuations, confirming stability as an equivalent condition for generalizing capability.
  • Probabilistic Bound: We establish an upper-bound on the generalization error in probability (Theorem 2) using the Sandwiched Rényi Divergence assuming an i.i.d. structure of the data and algorithm. Our unified framework leverages the non-commutative Hölder inequality to demonstrate an $O(1/\sqrt{n})$ convergence rate of the generalization error.
  • Lower Bound on Expected True Loss: We derive a reverse inequality linking expected true loss to empirical loss (Theorem 3). This establishes a tight relationship between empirical observations and true risk, recovering classical analogues [18] as limiting cases.
  • Stability of $(\varepsilon,\delta)$-QDP Learning Algorithms: We derive a mechanism-agnostic upper bound on the mutual information for $1$-neighbor $(\varepsilon,\delta)$-DP quantum learning algorithms (Theorem 4). Using grid-covering techniques, we generalize classical results [23] to show that privacy enforces stability with logarithmic sample size scaling.
  • Untrusted Processor Framework: We introduce Information-Theoretic Admissibility (ITA) for untrusted settings. We prove a fundamental quantum advantage: unlike classical settings where ITA precludes privacy, quantum non-commutativity allows for optimal algorithms that simultaneously maintain differential privacy.
@@ II Notations
Let $\mathcal{D}(\mathcal{H})$ denote the set of density operators on a finite-dimensional Hilbert space $\mathcal{H}$.
@@ III-A Learning Setup and Data Encoding
[…]
The Data Processor receives the classical-quantum input $\sum_{s}P_{S}(s)|s\rangle\langle s|\otimes\rho_{s}$ and executes a learning algorithm, modeled as a collection of quantum instruments $\mathcal{N}:=\{\mathcal{N}^{(s)}:\mathcal{H}^{\bm{\mathfrak{Tr}}}\to\mathcal{H}^{B}\}_{s\in\mathcal{S}}$. The output system $B\equiv WB^{\prime}$ comprises a classical hypothesis $W$ and a quantum residue $B^{\prime}$. The resulting joint state is given by:
@@ III-C Stability Implies Generalizability For Quantum Learning Algorithms
Definition 2 (Expected Empirical Loss [8, Definition $11$]).
@@
$\hat{L}_{\rho}(\mathcal{N})$ $:=\mathbb{E}_{(S,W)\sim{P}^{\mathcal{N}}_{SW}}[\mathrm{Tr}[L(S,W)(\sigma^{\mathcal{N}}_{S,W})^{\bm{\mathfrak{Te}}B^{\prime}}]]=\mathrm{Tr}[L^{S\bm{\mathfrak{Te}}WB^{\prime}}\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}].$
Definition 3 (Expected True Loss [26, Definition $19$]).
@@
$L_{\rho}(\mathcal{N})$ $:=\mathbb{E}_{(\overline{S},\overline{W})\sim P_{S}\times{P}^{\mathcal{N}}_{{W}}}\left[\mathrm{Tr}\left[L(\overline{S},\overline{W})\left(\rho^{\bm{\mathfrak{Te}}}_{\overline{S}}\otimes(\sigma^{\mathcal{N}}_{\overline{W}})^{B^{\prime}}\right)\right]\right]=\mathrm{Tr}[L^{S\bm{\mathfrak{Te}}WB^{\prime}}(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})].$
@@
We adopt the definition of expected true loss as proposed in [26] and not that of [8]. The authors in [26] give a rigorous justification for Definition 3 and argue that the definition proposed by [8, Definition $12$] is not a correct definition for the expected true loss.
@@
$\overline{\text{{gen}}}_{\rho}(\mathcal{N})$ $:=\left|\hat{L}_{\rho}(\mathcal{N})-L_{\rho}(\mathcal{N}))\right|=\left|\mathrm{Tr}[L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}]-\mathrm{Tr}\left[L^{S\bm{\mathfrak{Te}}WB}(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\right]\right|.$
@@
Definition 5 ( Classical-Quantum $\alpha$-Sub-Gaussianity).
For a fixed parameter $\alpha\in(0,\infty),$ the collection $\{L(w,s)\}$ of loss operators is said to be an $\alpha$-sub-Gaussian collection with respect to $\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}:=\sum_{(s,w)\in\mathcal{S}\times\mathcal{W}}P_{S}(s)|s\rangle\langle s|\otimes P^{\mathcal{N}}_{W}(w\mid s)|w\rangle\langle w|\otimes\rho^{\bm{\mathfrak{Te}}}_{s}\otimes(\sigma^{\mathcal{N}}_{w})^{B^{\prime}},$ if for every $\lambda\in\mathbb{R},$ it satisfies,
@@
Definition 5 naturally generalizes classical sub-Gaussianity. In the limit of trivial quantum systems ($\dim(\bm{\mathfrak{Te}})=\dim(B^{\prime})=1$), the operators $\sigma^{\mathcal{N}}_{W}$ and $\mathbb{I}^{B^{\prime}}$ become scalars, reducing $L(S,W)$ to a classical random loss function. Consequently, condition (6) collapses to the standard classical sub-Gaussian inequality $\mathbb{E}_{(S,W)}[e^{\lambda(L(S,W)-\mathbb{E}[L(S,W)])}]\leq e^{\frac{\lambda^{2}\alpha^{2}}{2}}$ with respect to $P_{S}\times{P}^{\mathcal{N}}_{{W}}$. We now present a theorem bounding the expected generalization error for quantum learning algorithms.
@@
See Appendix A for the proof.
The following corollary of Theorem 1, together with Definition 1, indicates that when the Data Processor employs a stable algorithm, the generalization error remains tightly bounded.
@@
While bounds on the expected generalization error provide a measure of average performance, robust learning requires guarantees that hold with high confidence for individual realizations of the algorithm. To address this, we prove a quantum version of [13, Corollary $2$] in term of the sandwiched Rényi divergence, derived under the assumption of i.i.d. data and loss observable decompositions.
Theorem 2.
Let $\mathcal{N}$ be a quantum learning algorithm. Assume that the associated collection of loss operators $\{L(s,w)\}$ satisfies the Conditional Classical-Quantum Local Sub-Gaussian condition (Definition 8). For any Sandwiched Rényi divergence order $\gamma>1$ and confidence level $\delta\in(0,1)$, the generalization error is bounded with probability at least $1-\delta$ as:
$\Pr_{(S,W)\sim P^{\mathcal{N}}_{SW}}\left\{\textnormal{gen}_{\rho}(\mathcal{N},S,W)\leq\sqrt{\frac{2\alpha^{2}}{n}\left(\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})+\frac{\gamma}{\gamma-1}\ln\frac{2}{\delta}\right)}\right\}\geq 1-\delta,$
where $\textnormal{gen}_{\rho}(\mathcal{N},S,W)$ is the generalization error random variable (Definition 9) and $\tilde{D}_{\gamma}$ denotes the Sandwiched Rényi divergence (defined in (19)).
Proof.
See Appendix C for the complete derivation, which includes a detailed discussion on the i.i.d. structure of the algorithm and the decomposition of the loss operators.
Complementing these upper bounds in expectation and probablity, we provide the following lower bound on the expected true loss in terms of the expected empirical loss.
Theorem 3.
Let $\mathcal{N}$ be a quantum learning algorithm with loss operators satisfying the Classical-Quantum Sub-Gaussian property (Definition 5) with parameter $\alpha>0$. For any sandwiched Rényi divergence order $\gamma>1$, the expected true loss is lower bounded by the empirical loss in the following exponential form,
$\exp\left(L_{\rho}(\mathcal{N})\right)\geq\hat{L}_{\rho}(\mathcal{N})\exp\left(-\left[\frac{\gamma\alpha^{2}}{2(\gamma-1)}+\frac{\gamma-1}{\gamma}\tilde{D}_{\gamma}(\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}\|\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\right]\right).$ (9)
Proof.
See Appendix B for the proof.
In Appendix D, we compare our upper-bounds on generalization error (Theorem 1 and Theorem 2) with prior works.